Hello. I am really stuck on my homework and need help. I am trying to solve this quadratic equation by factoring. Can anyone help? Thank you!
x^2+4x=12

Answer by ewatrrr(24785)


Hi
solving the following quadratic equation by factoring
x^2+4x = 12
x^2+4x -12 = 0 |must set to zero prior to factoring
factoring
(x ± __)(x ± __)|Ask ourselves what 2 numbers have a product of 12 and difference or 4
(x+6)(x-2)=0 Note:SUM of the inner product(6x) and the outer product(-2x) = 4x
(x+6)=0
x = -6
(x-2)=0
x = 2